Grade 12th passPhysical Chemistry

100 ml of an unknown solution of HCl is titrated with 50 ml 2M NaOH solution but after the addition the resultant solution required another 20 ml of 1N H2SO4 solution for Complete Neutralisation ? Find normality of Hcl solution ?

Deepak
9 Years agoGrade 12th pass

1 Answer

Deepak
9 Years ago
Equivalents of acid = Equivalents of base Eq of HCl + Eq of H2SO4 = Eq of NaOH100*N + 1*20 = 50*20*1100*N= 80N = .8